Find the x and y intercepts for 2x+3y=16
Answer by checkley75(3666)

SET X=0 & SOLVE FOR Y THEN SET Y=0 & SOLVE FOR X THUS:
2*0+3Y=16
3Y=16
Y=16/3 ANSWER FOR THE Y INTERCEPT (0,16/3)
2X+3*0=16
2X=16
X=16/2
X=8 ANSWER FOR THE X INTERCEPT (8,0)